# Scalable Two-Stage Stochastic Optimal Power Flow via Separable Approximation

### v1

@@ II-D Projection Algorithm
[…]
@@
[…]
@@
[…]
$m_{l}=n_{l}+\beta_{l-1}-\beta_{l}\quad\forall\ l=\underline{L}+1,\underline{L}+2,\dots,\overline{L}$ (20)
[…]
$\beta_{l}(m_{l}-m_{l+1})=0\quad\forall\ l=\underline{L},\underline{L}+1,\dots,\overline{L}$ (21)
[…]
@@
[…]
@@
    and assign, $m_{l}^{(k+1)}=d\quad\forall\ l=\eta,~{}\eta+1,\dots,l^{(k)}$, where $d$ is obtained from (25) with $p=\eta\ \text{and}\ q=l^{(k)}$. If no $\eta$ is found, assign $\eta=\underline{L}$.
@@
[…]
@@
    and assign, $m_{l}^{(k+1)}=d\quad\forall\ l=l^{k},l^{k}+1,\dots,\eta$ where $d$ is obtained from (25) with $p=l^{k}\ \text{and}\ q=\eta$. If no $\eta$ is found, assign $\eta=\overline{L}$. The Lagrange multipliers corresponding to the above solutions exactly matches with that of (20) and (21[15].
@@ II-E Statistical method for lower and upper bound
Since the proposed framework uses sample subgradient information to approximate the function, we provide a statistical method for evaluating the quality of the obtained solution through a posteriori analysis, utilizing the information available during the process of solving the stochastic optimization problem. We denote the objective function $H(x,y)+\mathbb{E}_{\xi}[Q(x,\xi)]$ in (1) as $F(x,y)$ and let $F^{*}$ be its optimal value. First, we provide the statistical upper bound, i.e., an estimate of the upper bound and its confidence interval at a chosen level. Let $(\hat{x},\hat{y})$ be the decisions obtained after solving (1) where the function is approximated using SPAR-OPF taking batches of $N^{\prime}$ samples from scenarios set $\Xi$. Consider that $p_{\xi}$ is the probability of scenario $\xi$ and $h_{|\Xi|}$ is the value of (1) where expected function is calculated by simulating all scenarios independently in the second stage $\textbf{P}_{\textbf{g2}}$ fixing the first stage decision $(\hat{x},\hat{y})$ as given in (26). This gives the upper bound for $F^{*}$. Taking expectation in the left side of (26) also gives the upper bound of $F^{*}$ as given in (27) where (1) is solved $M$ times taking batches of scenarios. The unbiased estimator, $\bar{h}_{|\Xi|}$, of $\mathbb{E}$($h_{|\Xi|}$) and variance $\sigma_{M}^{2}$ are given in (28) and (29), respectively. The confidence interval for the upper bound $UB_{M}$ with an approximated (1-$\alpha$) confidence is given in (30), which is justified by the central limit theorem with critical value $z_{\alpha}=f^{-1}(1-\alpha)$ where $f(z)$ is the cumulative distribution function of the standard normal distribution. Note that if the value of $M$ is low, the $t$-distribution should be used.
@@
The lower bound for $F^{*}$ can be obtained by leveraging the sample average approximation. The sample average approximation of $F_{N^{{}^{\prime}}}(x,y)$ with $N^{{}^{\prime}}$ samples is an unbiased estimator of $F(x,y)$ as given in left equality in (31). If $p_{N^{{}^{\prime}}}$ is an optimal value of (1) for $N^{{}^{\prime}}$ samples, then $\mathbb{E}(p_{N^{{}^{\prime}}})$ is a lower bound for $F(x,y)$ as given in (31). Taking minimum of $F(x,y)$ and using (31), we obtain (32). The unbiased estimator $\bar{p}_{N^{{}^{\prime}}}$ of $\mathbb{E}(p_{N^{{}^{\prime}}})$ and the variance are given in ( 33) and (34), respectively, where problem (1) is solved $M$ times taking $N^{{}^{\prime}}$ samples. The confidence interval for the Lower bound $LB_{N^{{}^{\prime}}}$ of $F^{*}$ with an approximated (1-$\alpha$) confidence is given by (35).
[…]
@@ III-A First Stage
[…]
@@ III-B Second Stage
$-z_{i}^{\phi,\xi}\leq\ v_{i}^{\phi,\xi}-v_{i,ref}^{\phi}\leq z_{i}^{\phi,\xi}\quad\forall\ i\in\mathcal{B},\quad\forall\ \phi\in\Phi$ (42)
@@
[…]
$\begin{gathered}U_{i}^{DG,\phi,\xi}\epsilon_{i,PV}^{\xi}-P_{i,load}^{\phi}\epsilon_{i,load}^{\xi}=\sum_{j\in\mathcal{C}_{i}}P_{ij}^{\phi,\xi}-\sum_{j\in\mathcal{P}_{i}}P_{ji}^{\phi,\xi}\\ \quad\forall\ i\in\mathcal{B}\backslash\{i_{sub}\},\quad\forall\ \phi\in\Phi\end{gathered}$ (44)
$\begin{gathered}U_{i}^{\phi,\xi}\epsilon_{i,PV}^{\xi}\tan\theta^{DG}-Q_{i,load}^{\phi,\xi}\epsilon_{i,load}^{\xi}=\sum_{j\in\mathcal{C}_{i}}Q_{ij}^{\phi,\xi}-\\ \sum_{j\in\mathcal{P}_{i}}Q_{ji}^{\phi,\xi}\quad\forall\ i\in\mathcal{B}\backslash\{i_{sub}\},\quad\forall\ \phi\in\Phi\end{gathered}$ (45)
@@
[…]
$\tilde{\textbf{r}}_{ij}={\rm Real}\{\alpha\alpha^{H}\}\otimes\textbf{r}_{ij}+{\rm Im}\{\alpha\alpha^{H}\}\otimes\textbf{x}_{ij}$
[…]
@@
Constraint (48) ensures that power flowing through line $ij$ is within its thermal rating $\textbf{S}^{rated}_{ij}$. This quadratic constraint is substituted with linear constraints (49), (50) and (51) using polygon based approximations [19].
[…]
$\textbf{P}_{ij}^{\xi^{2}}+\textbf{Q}_{ij}^{\xi^{2}}\leq(\textbf{S}_{ij}^{rated})^{2}\quad\forall\ ij\in\mathcal{E}$ (48)
[…]
$-\sqrt{3}(\textbf{P}_{ij}^{\xi}+\textbf{S}_{ij})\leq\textbf{Q}_{ij}^{\xi}\leq-\sqrt{3}(\textbf{P}_{ij}^{\xi}-\textbf{S}_{ij})\quad\forall\ ij\in\mathcal{E}$ (49)
[…]
$-\frac{\sqrt{3}}{2}\textbf{S}_{ij}\leq\textbf{Q}_{ij}^{\xi}\leq\frac{\sqrt{3}}{2}\textbf{S}_{ij}\quad\forall\ ij\in\mathcal{E}$ (50)
[…]
@@
[…]
where,
[…]
$\textbf{S}_{ij}=\textbf{S}_{ij}^{rated}\sqrt{\frac{2\pi/6}{\sin(2\pi/6)}}$
@@ III-C Load and PV Uncertainty Model
1:  Input: $\hat{\Xi},N_{t}$
2:  Output: $\Xi=\{\xi_{1},\xi_{2},...,\xi_{N_{t}}\}$
3:  $N_{D}^{\rm agg}\leftarrow\frac{|\hat{\Xi}|}{N_{t}}$
4:  initialize $s=1,b=1,e=N_{D}^{\rm agg}$
5:  while $e\leq|\hat{\Xi}|$ do
6:     for $h=1:24$ do
7:        $\alpha^{\rm load}(s)=\mathbb{E}[\alpha_{h}^{\rm load}(d)]_{d=b}^{e}$$\alpha^{PV}(s)=\mathbb{E}[\alpha_{h}^{PV}(d)]_{d=b}^{e}$
@@
The algorithm for scenario generation and reduction is presented in Algorithm 2, which takes $\hat{\Xi}$ scenarios consisting of hourly resolution historical normalized load and PV output and provides $N_{t}$ output scenarios. The scenario index of scenario $\xi_{s}\in\Xi$ is denoted by $s$ in the algorithm. The number of days to be aggregated $N_{D}^{\rm agg}$ in a stratum is calculated based on the total historical data available and the number of scenarios desired $N_{t}$. Then, aggregation is done for each hour for these aggregated days, the index of which starts from $b$ to $e$, to obtain scenarios. $\alpha_{h}^{\rm load}(d)\in\hat{\Xi}$ and $\alpha_{h}^{PV}(d)\in\hat{\Xi}$ denote the normalized load and PV multiplier at hour $h$ of day $d$, respectively. $\mathbb{E}[\alpha_{h}^{\rm load}(d)]_{d=b}^{e}$ and $\mathbb{E}[\alpha_{h}^{PV}(d)]_{d=b}^{e}$ denote the expectation of load and PV multiplier, respectively, corresponding to hour $h$ considering days indexed from $b$ to $e$. Then, once all 24 hours are done, the day’s index will be sided by $N_{D}^{\rm agg}$ to obtain scenarios considering the next strata of aggregated days and so on until the last data point in $\hat{\Xi}$ is reached.
@@ III-D SPAR-OPF algorithm for DG siting and sizing
Algorithm 3 shows the overall algorithm using SPAR-OPF for DG siting and sizing problem which takes the power flow model, scenarios set, and other required parameters222parameters that are used in modeling, see Section III-A and III-B as input and outputs the planning decisions: DG size $U_{i}^{DG^{\text{final}}}$ and siting decisions $\delta^{DG^{\text{final}}}$. The number of coordinate functions to be learned equals the number of buses, as DG decisions for each bus are linked to the second stage and we consider all buses as potential locations. Consider L as the maximum number of DG units that can be installed in a bus. Let $m_{i}^{l,(k)}\in m_{i}^{(k)}$ denote the slope of the coordinate function $g_{i}$ at point $l\in\mathcal{L}=[0,L]$ during iteration $k$ where $m_{i}^{(k)}$ is its slope vector. The slope $m_{i}^{l,(1)}$ of all coordinate functions is initialized to zero. The maximum number of iterations is set as $k^{\max}$. Then, at iteration $k$, with $m_{i}^{(k)}$ and using (12), the value of $g_{i}$ is updated. $g_{i}^{l,(k)}$ denotes its value at $l$ in the $k^{th}$ iteration. The first stage problem is then solved using either $\textbf{L}_{1}$ or $\textbf{E}_{1}$ and DG decisions (linking variables) $U_{i}^{DG^{(k)}}$ are passed to the second stage. Scenario $\xi$ is selected randomly from the scenario set $\Xi$ and uncertain parameters specified in (44) and (45) in the second stage problem III-B are updated. Then, the second stage problem is solved, and the dual information $\gamma_{i}^{(k)}$ of the linking constraints (43) is obtained. The slope of each coordinate function at the given first stage decision $U_{i}^{DG^{(k)}}$ is then adjusted using $\gamma_{i}^{(k)}$ where updated step size $\alpha^{(k)}$ is utilized. Using the projection algorithm as described in Section II-D, updated slope vector $n_{i}^{(k)}$ is then projected to the convex set to obtain $m_{i}^{(k+1)}$, which is utilized in the next iteration. The process continuous until the convergence criteria are met or the maximum iteration limit is reached.
@@
3:  $m_{i}^{l,(1)}\leftarrow 0\quad\forall\ i\in\mathcal{B},\ \forall\ l\in\mathcal{L}$
@@
9:        update $\epsilon_{i,PV}^{\xi},\epsilon_{i,load}^{\xi}$ in (44) and (45)
@@
14:     Update $\alpha^{(k)}$
15:     for each $\text{each bus}\ i\in\mathcal{B}$ do
16:        for $l=0,1,2,\dots,L$ do
17:           if $l=U_{i}^{DG^{(k)}}$ then
18:              $n_{i}^{l,(k)}=(1-\alpha^{(k)})m_{i}^{l,(k)}+\alpha^{(k)}\gamma_{i}^{(k)}$
19:           else
20:              $n_{i}^{l,(k)}=m_{i}^{l,(k)}$
21:           end if
@@
26:  Check the termination criteria
27:  Return: $U_{i}^{DG^{\text{final}}}\leftarrow U_{i}^{DG^{(k)}},\delta_{i}^{DG^{\text{final}}}\leftarrow\delta_{i}^{DG^{(k)}}$
@@ IV Results and Discussions
The default learning step size rule for SPAR-OPF, which we denote here as “step rule 1” is taken as $\alpha^{(k)}=20/(20+k)$ as suggested by [15]. Furthermore, different step size rules are tested to measure the robustness of the framework with respect to step size selection. Specifically, we define “step size rule 2” as $\alpha^{(k)}=1/k$, and “step size rule 3” as $\alpha^{(k)}=min(1,20/k)$ to cover different varieties. $k^{\max}$ and $\epsilon-radius$ are set to 100 and $10^{-4}$ for termination. The results, including the effects of increasing number of scenarios and system size, are then presented along with statistical validation. Finally, the computational comparison between the two formulations $\textbf{L}_{1}$ and $\textbf{E}_{1}$ is made. In the remainder of this paper, $\textbf{D}_{2}^{\textbf{v}}$ refers to two-stage stochastic optimization problem with voltage deviation minimization as an objective in the second stage, while $\textbf{D}_{2}^{\textbf{p}}$ refers to the one with power loss minimization as an objective.

### v2

@@ II-D Projection Algorithm
In this section, we present a simple projection rule that meets optimality conditions while avoiding optimization, thereby significantly accelerating the algorithm [15]. The projection of updated slope vector $n$ on convex set $\mathcal{M}$ is:
@@
 $\text{\quad s.t. }\quad m_{l}-m_{l+1}\leq 0\quad\forall\ l\in\{\underline{L},\underline{L}+1,\dots,\overline{L}\}$ (20)
@@
$m_{l}=n_{l}+\beta_{l-1}-\beta_{l}\quad\forall\ l\in\{\underline{L}+1,\underline{L}+2,\dots,\overline{L}\}$ (21)
$\beta_{l}(m_{l}-m_{l+1})=0\quad\forall\ l\in\{\underline{L},\underline{L}+1,\dots,\overline{L}\}$ (22)
Consider coordinates $p$ and $q$ with $p\leq q$ such that,
@@
    Find largest $\eta$ with $\underline{L}<\eta\leq l^{(k)}$ such that,
@@
    and assign, $m_{l}^{(k+1)}=d\quad\forall\ l\in\{\eta,~\eta+1,\dots,l^{(k)}\}$, where $d$ is obtained from (26) with $p=\eta\ \text{and}\ q=l^{(k)}$. If no $\eta$ is found, assign $\eta=\underline{L}$.
@@
    Find smallest $\eta$ with with $l^{(k)}\leq\eta<\overline{L}$ such that,
@@
    and assign, $m_{l}^{(k+1)}=d\quad\forall\ l\in\{l^{k},l^{k}+1,\dots,\eta\}$ where $d$ is obtained from (26) with $p=l^{k}\ \text{and}\ q=\eta$. If no $\eta$ is found, assign $\eta=\overline{L}$. The Lagrange multipliers corresponding to the above solutions exactly matches with that of (21) and (22[15].
@@ II-E Statistical method for lower and upper bound
The lower bound for $F^{*}$ can be obtained by leveraging the sample average approximation. The sample average approximation of $F_{N^{{}^{\prime}}}(x,y)$ with $N^{{}^{\prime}}$ samples is an unbiased estimator of $F(x,y)$ as given in left equality in (32). If $p_{N^{{}^{\prime}}}$ is an optimal value of (1) for $N^{{}^{\prime}}$ samples, then $\mathbb{E}(p_{N^{{}^{\prime}}})$ is a lower bound for $F(x,y)$ as given in (32). Taking minimum of $F(x,y)$ and using (32), we obtain (33). The unbiased estimator $\bar{p}_{N^{{}^{\prime}}}$ of $\mathbb{E}(p_{N^{{}^{\prime}}})$ and the variance are given in ( 34) and (35), respectively, where problem (1) is solved $M$ times taking $N^{{}^{\prime}}$ samples. The confidence interval for the Lower bound $LB_{N^{{}^{\prime}}}$ of $F^{*}$ with an approximated (1-$\alpha$) confidence is given by (36).
@@ III-A First Stage
 $\text{\quad\quad s.t. }\underline{P}^{DG}\delta_{i}^{DG}\leq U_{i}^{DG}u^{DG}\leq\overline{P}^{DG}\delta_{i}^{DG}\quad\forall\ i\in\mathcal{B}$ (38)
@@ III-B Second Stage
$z_{i}^{\phi,\xi}\leq\ v_{i}^{\phi,\xi}-v_{i,ref}^{\phi}\leq z_{i}^{\phi,\xi}\quad\forall\ i\in\mathcal{B},\quad\forall\ \phi\in\Phi$ (43)
@@
$\begin{gathered}U_{i}^{DG,\phi,\xi}\epsilon_{i,PV}^{\xi}-P_{i,{\rm load}}^{\phi}\epsilon_{i,{\rm load}}^{\xi}=\sum_{j\in\mathcal{C}_{i}}P_{ij}^{\phi,\xi}-\sum_{j\in\mathcal{P}_{i}}P_{ji}^{\phi,\xi}\\ \quad\forall\ i\in\mathcal{B}\backslash\{i_{sub}\},\quad\forall\ \phi\in\Phi\end{gathered}$ (45)
@@
$\text{where, }~\tilde{\textbf{r}}_{ij}={\rm Real}\{\alpha\alpha^{H}\}\otimes\textbf{r}_{ij}+{\rm Im}\{\alpha\alpha^{H}\}\otimes\textbf{x}_{ij}\hskip 36.135pt$
@@
Constraint (49) ensures that power flowing through line $ij$ is within its thermal rating $\textbf{S}^{\rm rated}_{ij}$. This quadratic constraint is substituted with linear constraints (50), (51) and (52) using polygon based approximations [17].
$\textbf{P}_{ij}^{\xi^{2}}+\textbf{Q}_{ij}^{\xi^{2}}\leq(\textbf{S}_{ij}^{\rm rated})^{2}\quad\forall\ ij\in\mathcal{E}$ (49)
$\sqrt{3}(\textbf{P}_{ij}^{\xi}+\textbf{S}_{ij})\leq\textbf{Q}_{ij}^{\xi}\leq-\sqrt{3}(\textbf{P}_{ij}^{\xi}-\textbf{S}_{ij})\quad\forall\ ij\in\mathcal{E}$ (50)
$\frac{\sqrt{3}}{2}\textbf{S}_{ij}\leq\textbf{Q}_{ij}^{\xi}\leq\frac{\sqrt{3}}{2}\textbf{S}_{ij}\quad\forall\ ij\in\mathcal{E}$ (51)
@@
$\text{where }~\textbf{S}_{ij}=\textbf{S}_{ij}^{\rm rated}\sqrt{\frac{2\pi/6}{\sin(2\pi/6)}}\,.\hskip 21.68121pt$
@@ III-C Load and PV Uncertainty Model
1:Input: $\hat{\Xi},N_{t}\text{ and }\textbf{Output: }\Xi=\{\xi_{1},\xi_{2},...,\xi_{N_{t}}\}$
[…]
3: initialize $s=1,b=1,e=N_{D}^{\rm agg}$
[…]
5:  for $h=1:24$ do
6:   $\alpha^{\rm load}(s)=\mathbb{E}[\alpha_{h}^{\rm load}(d)]_{d=b}^{e}$
7:   $\alpha^{PV}(s)=\mathbb{E}[\alpha_{h}^{PV}(d)]_{d=b}^{e}$
@@
The algorithm for scenario generation and reduction is presented in Algorithm 2, which takes $\hat{\Xi}$ scenarios consisting of hourly resolution historical normalized load and PV output and provides $N_{t}$ output scenarios. The scenario index of scenario $\xi_{s}\in\Xi$ is denoted by $s$ in the algorithm. The number of days to be aggregated $N_{D}^{\rm agg}$ in a stratum is calculated based on total historical data available and the number of scenarios desired $N_{t}$. Then, aggregation is done for each hour for these aggregated days, the index of which starts from $b$ to $e$, to obtain scenarios. $\alpha_{h}^{\rm load}(d)\in\hat{\Xi}$ and $\alpha_{h}^{PV}(d)\in\hat{\Xi}$ denote normalized load and PV multiplier at hour $h$ of day $d$, respectively. $\mathbb{E}[\alpha_{h}^{\rm load}(d)]_{d=b}^{e}$ and $\mathbb{E}[\alpha_{h}^{PV}(d)]_{d=b}^{e}$ denote the expectation of load and PV multiplier, respectively, corresponding to hour $h$ considering days indexed from $b$ to $e$. Then, once all 24 hours are done, the day’s index will be sided by $N_{D}^{\rm agg}$ to obtain scenarios considering the next strata of aggregated days and so on until the last data point in $\hat{\Xi}$ is reached.
@@ III-D SPAR-OPF algorithm for DG siting and sizing
Algorithm 3 shows the overall algorithm using SPAR-OPF for DG siting and sizing problem which takes the power flow model, scenarios set, and other required parameters used in the first and second stage models as input and outputs the planning decisions: DG size $U_{i}^{DG^{\text{final}}}$ and siting decisions $\delta^{DG^{\text{final}}}$. The number of coordinate functions to be learned equals the number of buses, as DG decisions for each bus are linked to the second stage and we consider all buses as potential locations. Consider L as the maximum number of DG units that can be installed in a bus. Let $m_{i}^{l,(k)}\in m_{i}^{(k)}$ denote the slope of the coordinate function $g_{i}$ at point $l\in\mathcal{L}=[0,L]$ during iteration $k$ where $m_{i}^{(k)}$ is its slope vector. The slope $m_{i}^{l,(1)}$ of all coordinate functions is initialized to zero. The maximum number of iterations is set as $k^{\max}$. Then, at iteration $k$, with $m_{i}^{(k)}$ and using (12), the value of $g_{i}$ is updated. $g_{i}^{l,(k)}$ denotes its value at $l$ in the $k^{th}$ iteration. The first stage problem is then solved using either $\textbf{L}_{1}$ or $\textbf{E}_{1}$ and DG decisions (linking variables) $U_{i}^{DG^{(k)}}$ are passed to the second stage. Scenario $\xi$ is selected randomly from the scenario set $\Xi$ and uncertain parameters specified in (45) and (46) in the second stage problem III-B are updated. Then, the second stage problem is solved, and if it is feasible, the dual information $\gamma_{i}^{(k)}$ of the linking constraints (44) is obtained. The slope of each coordinate function at the given first stage decision $U_{i}^{DG^{(k)}}$ is then adjusted using $\gamma_{i}^{(k)}$ where updated step size $\alpha^{(k)}$ is utilized. Using the projection algorithm as described in Section II-D, updated slope vector $n_{i}^{(k)}$ is then projected to the convex set to obtain $m_{i}^{(k+1)}$, which is utilized in the next iteration. On the other hand, if the second stage is infeasible, a feasibility cut of the form (18) is added to the first-stage problem. The process continues until the termination criterion is met.
@@
[…]
@@
9:   update $\epsilon_{i,PV}^{\xi},\epsilon_{i,{\rm load}}^{\xi}$ in (45) and (46)
@@
15:   Update $\alpha^{(k)}$
16:   for $\text{each bus}\ i\in\mathcal{B}$ do
17:    for $l=0,1,2,\dots,L$ do
18:     if $l=U_{i}^{DG^{(k)}}$ then
19:      $n_{i}^{l,(k)}=(1-\alpha^{(k)})m_{i}^{l,(k)}+\alpha^{(k)}\gamma_{i}^{(k)}$
20:     else
21:      $n_{i}^{l,(k)}=m_{i}^{l,(k)}$
22:     end if
@@
26:  else
27:   Add feasibility cut (18) to first stage problem (III-A)
28:  end if
29:end for
30: Check the termination criteria
31:Return: $U_{i}^{DG^{\text{final}}}\leftarrow U_{i}^{DG^{(k)}},\delta_{i}^{DG^{\text{final}}}\leftarrow\delta_{i}^{DG^{(k)}}$
@@ IV Results and Discussions
The default learning step size rule for SPAR-OPF, which we denote here as “step rule 1” is taken as $\alpha^{(k)}=20/(20+k)$ as suggested by [15]. Furthermore, different step size rules are tested to measure the robustness of the framework with respect to step size selection. Specifically, we define “step size rule 2” as $\alpha^{(k)}=1/k$, and “step size rule 3” as $\alpha^{(k)}=min(1,20/k)$ to cover different varieties. $k^{\max}$ and $\epsilon-radius$ are set to 100 and $10^{-4}$. The results, including the effects of increasing number of scenarios and system size, are then presented along with statistical validation. Finally, the computational comparison between the two formulations $\textbf{L}_{1}$ and $\textbf{E}_{1}$ is made. In the remainder of this paper, $\textbf{D}_{2}^{\textbf{v}}$ refers to two-stage stochastic optimization problem with voltage deviation minimization as an objective in the second stage, while $\textbf{D}_{2}^{\textbf{p}}$ refers to the one with power loss minimization as an objective.